# Classification of Breast Cancer Histopathology Images using a Modified Supervised Contrastive Learning Method

### v1

@@ 4.3 Results
In this section, we compare the results of the proposed method with those of the methods, PFTAS [22], pdMISVM [21], MPCS-OP [6], MPCS-RP [6], DenseNet121 [17], and ResHist-Aug [9]. Table 1 and Table 2 illustrate the image-level and patient-level accuracy achieved by different methods, respectively.
As shown in Table 1, the model proposed in this study has outperformed the existing methods in terms of accuracy across three magnification factors while exhibiting a slight decrease in the other magnification factor. On average, this method has demonstrated an improvement of $1.45\%$ in the image level and $1.42\%$ in the patient level accuracy compared to the previous models.
The results from previous methods demonstrate that as the magnification factor increases, the accuracy decreases. This trend can be attributed to higher magnification images capturing more intricate and internal details of the cells, in contrast to lower magnification images that primarily capture the overall cell structure. Consequently, the higher magnification images pose a more significant challenge for neural networks to learn and classify accurately. However, this is not the case in the proposed model, as we were able to achieve a high accuracy in all four magnification factors. This can be attributed to the characteristic of contrastive methods, where learning is derived directly from the underlying data and not only the labels.
@@
To evaluate the generalization capability of the trained model, we utilized the weights of models that were previously trained on each of the four magnification factors of the BreakHis dataset. These weights served as the initial starting point for training the BACH dataset. As shown in Table 3, the model can achieve an accuracy above 90% after only 20 epochs, indicating that this method has successfully learned discriminative features specific to histopathology images and can generalize well to other histopathology datasets with minimal additional training. Moreover, a decreasing trend in the accuracy can be observed as the magnification factor increases, which suggests that the BACH dataset images are closer to the 40X magnification factor of the BreakHis. This observation is further supported by visually inspecting the two datasets’ images.
@@
Method Comb1 Comb2 Comb3 Comb4 Comb5 Comb6 Comb7
Sup
Elim
Self
Accuracy 93.68 89.50 91.16 95.08 93.86 91.35 96.74
[…]
Table 4 analyses the impact of each of the loss functions 5, 6, and 7 and their combinations. These experiments have only been conducted on one fold, with the primary objective being the comparison of various loss functions rather than a comparison with previous methods. The results displayed in the table indicate that $\ell_{sup}$ has the most significant effect in improving the outcomes. This can be due to the fact that $\ell_{sup}$ leverages the labels to eliminate the impact of false negatives in the numerator. Moreover, this loss function aims to minimize the similarity between the anchor and both negative and positive pairs within a batch in the denominator. By doing so, $\ell_{sup}$ effectively prevents the network from being disrupted in the presence of a not similar positive pair. Another aspect of this loss function is the selection of positive pairs, which is done among all patients with similar diseases. By bringing the representations of images from different patients closer together, we enable the network to generalize its learning to the images of unseen patients in the test set.
The $\ell_{self}$ has the most negligible impact among all the losses. This loss is similar to $\ell_{sup}$, with the only difference being the positive pairs. Including this term increases the influence of the positive pair generated through data augmentation, leading to an improvement in accuracy.
@@ 5 Conclusion and Future Work
In the proposed method, a potential concern arises from considering all images with the same label as positive pairs during the initial representation stage. This can be problematic since cancerous images can exhibit significant variations, even within a single patient. In this approach, all images, including those from the same patient and different ones, are treated as positive pairs. Consequently, some pairs that were considered positive may not exhibit sufficient similarity and actually present a challenge for the network. In attempting to increase their similarities, the loss function can lead to a deviation from the model’s ultimate goal. To overcome this issue, our future work will aim to extract pairs more intelligently, beyond relying solely on the labels.
@@ 6.3 Author Contributions
All authors contributed to the study’s conception and design. Material preparation, data collection, and analysis were performed by all authors. The first author wrote the first draft of the manuscript, and all authors commented on several versions of the manuscript. All authors read and approved the final manuscript.

### v2

@@ 4.3 Results
In this section, we compare the results of the proposed method with those of the methods, PFTAS [28], pdMISVM [27], MPCS-OP [7], MPCS-RP [7], DenseNet121 [21], ResHist-Aug [11], MSB[30], MIB[31], GLNET[17], and quadtree16[37]. Table 1 and Table 2 illustrate the image-level and patient-level accuracy achieved by different methods, respectively.
As shown in Table 1, the model proposed in this study has outperformed the existing methods in terms of accuracy across three magnification factors while exhibiting a slight decrease in the other magnification factor. On average, this method has demonstrated an improvement of $1.45\%$ in the image level, and $0.31\%$ in the patient level accuracy compared to the previous models.
@@
To evaluate the generalization capability of the trained model, we utilized the weights of models that were previously trained on each of the four magnification factors of the BreakHis dataset. These weights served as the initial starting point for training the BACH dataset. As shown in Table 4, the model can achieve an accuracy above 90% after only 20 epochs, indicating that this method has successfully learned discriminative features specific to histopathology images, and can generalize well to other histopathology datasets with minimal additional training. Moreover, a decreasing trend in the accuracy can be observed as the magnification factor increases which suggests that the BACH dataset images are closer to the 40X magnification factor of the BreakHis. This observation is further supported by visually inspecting the two datasets images.
4.3.1 Ablation Study
The structure of our presented loss function is similar to Eq. 9. By introducing a coefficient $\lambda$, and the $Q$ set, we have incorporated the effect of the Eq. 10. Additionally, by utilizing augmentations for creating a positive pair for images, we have also integrated the impact of the Eq. 8 into our approach.
Method Comb1 Comb2 Comb3 Comb4 Comb5 Comb6 Comb7
Sup
Elim
Self
Accuracy 93.68 89.50 91.16 95.08 93.86 91.35 96.74
[…]
Table 3 analyses the impact of each of the loss functions 8, 9, and 10 and their combinations. These experiments have only been conducted on one fold, with the primary objective being the comparison of various loss functions.The results displayed in the table indicate that $\ell_{sup}$ has the most significant effect in improving the outcomes. This can be due to the fact that $\ell_{sup}$ leverages the labels to eliminate the impact of false negatives in the numerator. Moreover, this loss function aims to minimize the similarity between the anchor and both negative and positive pairs within a batch in the denominator. By doing so, $\ell_{sup}$ effectively prevents the network from being disrupted in the presence of a not similar positive pair. Another aspect of this loss function is the selection of positive pairs, which is done among all patients with similar diseases. By bringing the representations of images from different patients closer together, we enable the network to generalize its learning to the images of unseen patients in the test set.
The $\ell_{self}$ has the most negligible impact among all the losses. This loss is similar to $\ell_{sup}$, with the only difference being the positive pairs. Including this term increases the influence of the positive pair generated through data augmentation, which in turn leads to an improvement in accuracy.
5 Discussion
The results from previous methods demonstrate that as the magnification factor increases, the accuracy decreases. This trend can be attributed to higher magnification images capturing more intricate and internal details of the cells, in contrast to lower magnification images that primarily capture the cell structure and overall patterns. Consequently, the higher magnification images pose a more significant challenge for neural networks to learn and classify accurately. However, this is not the case in the proposed model as we were able to generally achieve a high accuracy in all of the four magnification factors. This can be attributed to the characteristic of the proposed loss function. The combination of self-supervised, supervised contrastive, and elimination ensures that the model learns a comprehensive set of features. Self-supervised learning helps capture a wide range of visual patterns, while supervised contrastive loss refines these features based on class labels. Elimination part of loss further ensures that only reliable features are considered. This holistic approach helps the model to effectively capture and utilize detailed features present in high magnification images as well as the low magnifications. This uniformly high performance across all magnification levels is the main reason our overall accuracy generally surpasses that of the models presented in previous cited experiments.
@@ 6 Conclusion and Future Work
In the proposed method, a potential concern arises from considering all images with the same label as positive pairs during the initial representation stage. This can be problematic since cancerous images can exhibit significant variations, even within a single patient. In this approach, all images, including those from the same patient and different ones, are treated as positive pairs. Consequently, some pairs which were considered positive may not exhibit sufficient similarity, and actually present a challenge for the network. The loss function, in attempting to increase their similarities, can lead to a deviation from the model’s ultimate goal. To overcome this issue, our future work will aim to extract pairs more intelligently, beyond relying solely on the labels.
@@ 7.3 Author Contributions
All authors contributed to the study conception and design. Material preparation, data collection, and analysis were performed by all authors. The first draft of the manuscript was written by the first auther and all authors commented on several versions of the manuscript. All authors read and approved the final manuscript.
@@ 7.4 Ethics Approval
Biography
Matina Mahdizadeh Sani
Matina Mahdizadeh Sani is currently a graduate student at University of Waterloo, Canada. She got her bachelor’s degree from Sharif university of technology, Tehran, Iran in 2023. Her field of study is machine learning and computer vision.
Ali Royat
Ali Royat is a graduate student (B.Sc. and M.Sc.) in Electrical Engineering at the Sharif University of Technology, Tehran, Iran. His expertise lies in the application of deep learning to Computer Vision and signal processing.
Mahdieh Soleymani Baghshah
Mahdieh Soleymani received her BS, MS, and PhD degrees from the Department of Computer Engineering, Sharif University of Technology, Iran, in 2003, 2005, and 2010, respectively. Her MS and PhD theses were in the field of Machine Learning. She joined Sharif University of Technology as an assistant professor of Computer Engineering and founded the Machine Learning Laboratory (MLL) in 2012. She is currently an associate professor at Sharif University of Technology and the director of MLL. Her research interests include machine learning and deep learning. She has worked on various types of data including images, texts, videos, omics, and financial time series.